What is the probability that the height of a child, picked at random, is less than 104 cm?


1
Expert's answer
2021-03-09T05:10:49-0500

Five hundred children participated in a field demonstration. Their heights averaged 110

cm with a standard deviation of 6 cm.

The probability that X<104 is equal to the blue area under the curve.



Since "\mu=110"  and "\sigma=6"  we have:

"P(X<104) = P(X-\mu<104-110)=P(\frac{X-\mu}{\sigma}<\frac{104-110}{6})"

Since 

"\frac{x-\mu}{\sigma} = Z"  and "\frac{104-110}{6} = -1"

 we have:

"P(X<104)=P(Z<-1)"

Use the standard normal table to conclude that:

"P(Z<-1)=0.1587"

Answer: 0.1587.